# MeshPad: Interactive Sketch-Conditioned Artist-Designed Mesh Generation and Editing

### v2

@@ Appendix E Additional Visual Results
Appendix F Perceptual Study Details
Figure 12: Screenshot of our perceptual study. Left: single-method evaluation. Right: dual-method comparison. Each participant is required to complete 10 single-method evaluations and 10 dual-method comparison tasks.
@@ Appendix F Perceptual Study Details
Table 5: User ratings (ranged $1$-$5$) on generation and editing results of our hand-drawn sketch evaluation set. We benchmark against LAS [63] and SENS [7], with outputs further refined using MeshAnythingV2 (MA) [17] as a post-processing baseline. For mesh generation, participants evaluate mesh quality (GQ) and sketch matching (GM). For mesh editing, participants rate edited mesh quality (EQ), edited sketch matching (EM), and edited mesh consistency (EC) of the unedited part.

### v3

@@ Appendix D Additional Qualitative Experiments
Mesh-to-Sketch Alignment. Our method achieves superior performance on sketch alignment. Fig. 12 shows that our method can (a, b): generate shapes aligning with the length and size of the sketch input; (c): edit the middle part of the mesh; (d): complete broken shapes via addition. While our method achieves better sketch alignment metrics than baseline methods, it does not guarantee exact alignment between generated meshes and sketches, since our model aims at high-fidelity shape generation from freehand sketches.
Figure 12: Qualitative examples showing the editing capability of our model. (a, b): sensitivity of edited shapes to the sketch input. (c): editing the middle part of the shape. (d): the added part (cyan) connects to the existing part.
Figure 13: Qualitative results on Objaverse [21] (a, b) and DeformingThings4D [37] (c, d). We fine-tune our networks on the two datasets and show their scalability to multiple categories.
@@ Appendix E Additional Visual Results
In Fig. 13, we show the mesh editing results of our networks fine-tuned on Objaverse [21] and DeformingThings4D [37]. For Objaverse fine-tuning, we filter the meshes with fewer than $768$ triangles to obtain around $9,000$ meshes. For DeformingThings4D, we sample around $1,000$ meshes from all animations and decimate the mesh to fit the limit of our setup. The result shows the scalability of our method to a large variety of meshes.
Method #Faces #Components Non-manifold Edges% #Self Intersections
LAS 33,573 1.0 0 0
LAS-MA 1,114 51.9 2.05 1,738
SENS 33,859 2.5 0 0
SENS-MA 1,513 73.4 2.22 1,787
Ours 231 12.3 2.01 259
Training Data 285 12.0 2.13 3,285
Table 5: Topological statistics of edited meshes vs. train data. Our method generates meshes that align best with the training data with significantly reduced self-intersections. Note that LAS and SENS generate over-tessellated meshes that deviate from the training data.
@@
Table 6: User ratings (ranged 11-55) on generation and editing results of our hand-drawn sketch evaluation set. We benchmark against LAS [65] and SENS [7], with outputs further refined using MeshAnythingV2 (MA) [17] as a post-processing baseline. For mesh generation, participants evaluate mesh quality (GQ) and sketch matching (GM). For mesh editing, participants rate edited mesh quality (EQ), edited sketch matching (EM), and edited mesh consistency (EC) of the unedited part.
Appendix F Topology of Generated Meshes
As shown in Tab. 5, our method generates meshes with similar topological statistics to the training data, while meshes of baseline methods deviate from it. LAS and SENS meshes are over-tessellated marching-cube meshes and are not ideal, even with no non-manifold edges and no self-intersections. We also surprisingly find that the number of self-intersections of our method is significantly less than the training data. We believe the reason is that the model learns to use the simplest way to represent a surface, while in ShapeNet, there are often redundant triangles.
Appendix G Perceptual Study Details
Figure 14: Screenshot of our perceptual study. Top: single-method evaluation. Bottom: dual-method comparison. Each participant is required to complete 10 single-method evaluations and 10 dual-method comparison tasks.